5 Consider a discrete random variable X with the probability mass function rp(x) Consider Y = g( X ) =- 0.2 0.4 0.3 0.1 a) Find the probability distribution of Y. b Find the expected value of Y, E(Y). Does μ Y equal to g(Hy )? 4
Answer #1

from above below is probability distribution of Y:

x y=1/x P(y)
4 1/4 0.10
3 1/3 0.30
2 1/2 0.40
1 1    0.20

b) E(Y) =\muY =\sumyP(y) =(1/4)*0.1+(1/3)*0.3+(1/2)*0.4+1*(0.2)=0.525

as  \mux =\sumxP(x) =1*0.2+2*0.4+3*0.3+4*0.1=2.3

therefore g( \mux)=1/2.3 =0.4348 which is not equal to \muY